OpenStudy (anonymous):

A basketball team has won 50 games of 75 played. The team still has 45 games to play. How many of the remaining games must the team win in order to win 60% of all games played during the season?

OpenStudy (anonymous):

how would I solve this?

OpenStudy (anonymous):

played = 75 still to be played = 45 so total games to be played in the season = 75 + 45 = 120 60% of total games = 70/100 * 120 = 8400/100 = 84 so they have to win 84 games in the season to have a win %age of 60% they hv already won 50 games so they need to win 84 - 50 = 34 more games